# Large Language Model Performance Benchmarking on Mobile Platforms: A Thorough Evaluation

### v1

@@ 4.1.2 Performance on GPUs
Unlike CPUs, which are all from Arm, GPUs across vendors may have different architectures. Thus, it is challenging for developers to optimize for different GPUs. Figure 3 illustrates the end-to-end latency of MLC LLM across four high-tier GPUs from Arm Mali and Qualcomm Snapdragon. Note that Huawei Matepad11 Pro and Huawei Nova7 are excluded from this comparison due to subpar performance on devices with low RAM (GB). Among the tested GPUs, Mali GPUs, particularly the Mali-G78, exhibit significantly slow prefill speeds. Consequently, we only consider the performance with 64-token prompts. For the Mali-G78, due to its extremely poor prefill performance, we only present results for 32-token prompts as a reference.
@@
The Adreno GPU consistently outperforms the Mali GPU in overall performance. Notably, Mali GPUs show unusually poor prefill performance. As for decoding, the Adreno delivers faster decoding speed than the Mali-G. Similarly, the Adreno achieves faster decoding speed compared to the Mali-G. Although Mali GPUs have poor prefill performance, Mali-G still achieves speedup in decoding than older-generation Adreno . Memory usage across the GPUs is nearly identical at around GB, with a slight increase to GB observed on the Vivo Pad3 Pro.
4.1.3 Performance on Specialized AI Accelerators
Recent years have seen the introduction of specialized AI accelerators in mobile SoCs, such as the Snapdragon Hexagon NPU, Kirin NPU, and MediaTek APU. However, these accelerators face several challenges, including limited support for dynamic shapes and floating-point operations, which complicates their use for LLM inference [46]. The Hexagon NPU, being the only mobile NPU with an open instruction set architecture, has garnered significant research interest. Successful deployments of mobile LLM models on the Snapdragon 8 Gen 3 have been reported. We summarize the findings from mllm, PowerInfer-, and Qualcomm AI Hub in Figure 4.
Qualcomm AI Hub, the official tool developed by the manufacturer, demonstrates a remarkable prefill speed of tokens per second, representing a -fold increase compared to both CPU and GPU. Similarly, mllm and PowerInfer-2 achieve substantial improvements in prefill speed, with an order of magnitude improvement in prefill speed. However, the decoding speed remains just slightly better than that of CPU and GPU based solutions. As we all know, prefill is compute-bound and decoding is memory-bound. This distinction likely accounts for the significant improvement in prefill speed observed with the NPU, whereas its decoding speed remains comparable to that of CPU/GPU.
4.1.4 Comparisons and Summary
1.
Given that mobile CPUs are built on the same Arm architecture and share common instruction sets, deploying LLMs on CPU is more straightforward and does not require specialized customization. As a result, CPUs have emerged as a preferred option for LLM inference on mobile platforms. However, our results show that they are limited in performance and thus not the optimal solution.
2.
@@ 4.1.4 Comparisons and Summary
GPUs, despite their theoretical advantages for LLMs due to their parallel float computing capabilities, do not show a clear advantage in practice. What’s counterintuitive is that MLC LLM on GPUs performs worse in prefill speed compared to llama.cpp on CPUs, a phenomenon that we will provide a detailed analysis in Section 5.1.3.
3.
In contrast, early trials with the Hexagon NPU demonstrate a significant advantage for specialized AI accelerators in inference tasks. However, due to their limited support for operators, adapting models and algorithms to specialized AI accelerators remains a challenging task for developers.
4.2 Battery Consumption
Table 2 compares the power consumption and inference speed between the Vivo Pad3 Pro and Huawei MatePad 12.6 Pro. To obtain the power consumption, we read the data provided by the OS settings, which records and displays the power consumption of a certain application. Old-generation devices are excluded from this comparison due to significant differences in battery capacity and power-saving policies compared to new-generation hardware. We perform consecutive rounds of inference on the CPU, with a prefill length of and a generation length of . The Dimensity 9300 exhibites a speedup in prefill and a speedup in decoding, while consuming only of the power used by the Kirin E. These results suggest that top-tier SoCs are capable of delivering substantial energy savings alongside improved performance.
Most modern smartphones feature battery capacities between mAh and mAh. At a power consumption rate of mAh per inference round, a device could theoretically support over rounds of inference. This aligns with the findings in [25], reinforcing the feasibility of LLM inference on mobile devices. However, a significant rise in device temperature during inference suggests that LLM inference is a power-intensive task. We recorded an increase from °C to °C during a single inference on Huawei Matepad12.6 Pro. Thus, optimizing the power efficiency of LLMs on mobile devices is crucial for sustainable and long-term deployment.
@@ 4.2 Battery Consumption
(mAh/round) | (token/s) | (token/s) | |
@@
5 Performance: Developers’ Perspectives
@@ 5.1 Hardware Capabilities
5.1.1 Multi-threads on CPU Cores
@@
Most popular mobile SoCs utilize the "big.LITTLE" architecture for their CPUs, which balances performance with power efficiency. This configuration typically includes multiple cores organized into two distinct clusters: "big" (prime and performance cores) and "little" (efficiency cores), as illustrated in Table 3. While it is commonly assumed that high-load tasks are best handled by the "big" cores for optimal performance, our tests reveal that the ideal core configuration can vary across the two stages of LLM inference.
@@
Figure 5 illustrates how these core configurations impact inference speed across the four tested devices. During the prefill stage, prompt processing speed is primarily driven by the big cores. The contribution of the little cores to acceleration depends on the performance gap between the big and little cores. On devices with limited computational power, additional efficiency cores can help accelerate inference. For instance, utilizing all cores delivers the best performance on the Kirin 9000E. However, for devices with more powerful big cores, adding little cores can actually reduce performance. For example, on the Snapdragon 8 Gen 3, incorporating two little cores results in a slowdown in inference speed. This highlights the importance of optimizing core configurations based on each device’s specific CPU capabilities to maximize efficiency. Interestingly, the Dimensity 9300 benefits from using all cores because of its All-Big-Core architecture [30].
In the decoding phase, adding more little cores typically leads to a decline in performance. For instance, both the Kirin 9000E and Snapdragon 8+ Gen 1 achieve their peak decoding speeds with four active big cores. This indicates that the maximum decoding speed on a device is largely dictated by the performance of the big cores. Unlike the prefill stage, decoding is more heavily constrained by memory bandwidth. This is evident in the fact that adding more big cores results in a smaller performance improvement compared to prefill, while incorporating little cores leads to a more noticeable performance drop.
5.1.2 Speedup with Special Machine Instructions
@@
Figure 6 highlights a notable performance boost. When focusing on peak inference performance, prefill speed is accelerated by under the same configuration. This speedup is primarily achieved by increasing throughput, allowing more computations to be processed within a single instruction. On the Snapdragon 8 Gen 3 (Xiaomi 14 Pro), adding all efficiency cores (little cores) results in a decline compared to using only the big cores. This suggests that the lower frequency and IPC (instructions per cycle) of efficiency cores may hinder performance during high-demand tasks.
@@
5.1.3 Utilization of GPU units
We note an intriguing exception regarding performance on Mali GPUs. The Mali-G Immortalis MP12, found in the Dimensity 9300, has better hardware parameters but has poorer performance than Adreno in Snapdragon SoCs. We measure the maximum throughput for FP16 operations on each device using clpeak[24]. It is a benchmarking tool intended for calculating GPU performance by executing different OpenCL kernels on a particular device. We also measure the actual maximum memory bandwidth using clpeak and present the maximum theoretical bandwidth in Table 4. From Table 4, we can see that the Mali-G720 Immortalis MP12 exhibits times the throughput of the Adreno in float16 operations. However, in Figure 3, it shows a prefill speed that is times slower.
The possible reason for this discrepancy comes from the GPU utilization. Figure 7 presents the average utilization and memory bandwidth during inference on Mali-G720 and Adreno 750. It shows that MLC LLM does not efficiently utilize the Mali GPU ALUs. During the compute-bound prefill stage, the Mali-G720’s arithmetic unit utilization averages less than , while the Adreno achieves around . Similarly, in the memory-bound decoding stage, the Mali-G achieves a real-time memory read bandwidth of GB/s, which is only roughly two-thirds of the GB/s of the Adreno 750. These results indicate existing LLM implementations cannot fully utilize the available computing resources on mobile GPUs. It implies that while OpenCL serves as a general-purpose backend, it is essential to implement GPU-specific kernels to fully leverage the hardware capabilities.
5.2 Impact of DVFS

### v3

@@ IV-A2 Performance on GPUs
Unlike CPUs, which are all from Arm, GPUs across vendors may have different architectures. Thus, it is challenging for developers to optimize for different GPUs. Figure 3 illustrates the end-to-end latency of MLC LLM across five high-tier GPUs from Arm Mali, Qualcomm Snapdragon, and Apple. It should be noted that Huawei MatePad 11 Pro and Huawei Nova 7 exhibit frequent crashes when running MLC LLM, which prevents the collection of reliable measurement results. Consequently, these two devices are excluded from the comparison. Among the tested GPUs, Mali GPUs, particularly the Mali-G78, exhibit significantly slow prefill speeds. Consequently, we only consider the performance with 64-token prompts. For the Mali-G78, due to its extremely poor prefill performance, we only present results for 32-token prompts as a reference.
@@
Although the decode speed of the Apple M1 is comparable to that of the Adreno 750, its prefill speed is nearly higher. This result highlights the huge performance gap between Apple SoCs and non-Apple SoCs. The Adreno GPU consistently outperforms the Mali GPU in overall performance. Notably, Mali GPUs show nearly unusable prefill performance. When comparing products from the similar tier, the Adreno delivers faster decoding speed than the Mali-G. Similarly, the Adreno achieves faster decoding speed compared to the Mali-G. Although Mali GPUs have poor prefill performance, Mali-G still achieves speedup in decoding than older-generation Adreno . To look into this phenomenon, we will provide a detailed analysis in Section V-B. Memory usage across the GPUs is nearly identical at around GB, with a slight increase to GB observed on the Vivo Pad3 Pro.
Implication 2
@@
For non-Apple GPUs, despite their theoretical advantages for large language models (LLMs) due to their parallel floating-point computing capabilities, do not exhibit a clear advantage in practice. Counterintuitively, MLC LLM on GPUs performs worse in prefill speed compared to Llama.cpp on CPUs. The reason could be that optimizing LLMs on mobile GPUs is more complex than on CPUs due to the diversity of GPU architectures.
IV-A3 Performance on Specialized AI Accelerators
Specialized AI accelerators have recently been integrated into mobile SoCs, such as the Snapdragon Hexagon NPU, Kirin NPU, and MediaTek APU. However, these accelerators face several challenges, including limited support for dynamic shapes and floating-point operations, which complicates use for LLM inference [35]. The Hexagon NPU, being the only mobile NPU with an open instruction set architecture, has garnered significant research interest. Successful deployments of mobile LLM models on the Snapdragon 8 Gen 3 have been reported. We collect the throughput reported in [38](mllm) and PowerInfer- [35], and compare with the performance on CPU and GPU.
MLLM and PowerInfer- achieve roughly an order-of-magnitude improvement in prefill speed over CPU, with MLLM reaching tokens/s for a -token prompt (versus tokens/s on CPU) and PowerInfer- achieving tokens/s for a -token prompt (versus tokens/s on CPU). Given the memory-bound nature of the decode stage, a hybrid inference scheme—using the NPU for prefill and the CPU or GPU for decode—may be a suitable approach for future optimization.
Implication 3
Although manufacturers provide limited support for NPUs, early trials with the Hexagon NPU have shown a substantial advantage in inference tasks. In order to drive this forward, more efforts should be devoted to adapting models and algorithms to specialized AI accelerators.
@@ IV-A3 Performance on Specialized AI Accelerators
| (Prompt, Generation) = (64, 128)Tokens | (mAh/round) | (token/s) | (token/s) |
@@
IV-B Battery Consumption
Table II compares the power consumption and inference speed between the Vivo Pad3 Pro and Huawei MatePad 12.6 Pro. To obtain the power consumption, we read the data provided by the OS settings, which record and display the power consumption of a certain application. Old-generation devices are excluded from this comparison due to significant differences in battery capacity and power-saving policies compared to new-generation hardware. We conduct consecutive rounds of inference on the CPU using threads to balance performance and energy consumption, with a prefill length of and a generation length of . Compared with the Kirin E (Huawei Matepad12.6 Pro), the Dimensity 9300 (Vivo Pad3 Pro) exhibits a speedup in prefill and a speedup in decoding, while consuming only of the power. These results suggest that top-tier SoCs are capable of delivering substantial energy savings alongside improved performance.
In recent years, most smartphones have been equipped with battery capacities ranging from mAh to mAh. At a power consumption rate of mAh per inference round, a device could theoretically support over rounds of inference. This result aligns with the findings in [8], reinforcing the feasibility of LLM inference on mobile devices. However, a significant rise in device temperature during inference suggests that LLM inference is a power-intensive task. We recorded an increase from °C to °C during a single inference on Huawei Matepad12.6 Pro. Thus, optimizing the power efficiency of LLMs on mobile devices is crucial for sustainable and long-term deployment.
IV-C Parallel with Background Tasks (non-AI tasks)
In real-world scenarios, users often run inference alongside other processes, which can create a gap between peak and actual performance. To gain a deeper understanding of performance under typical conditions, we expand our evaluation of LLM inference by incorporating concurrent background workloads.
The experimental configuration remains unchanged from the one described in Section 1, with the exception that a music application is launched after rebooting the device and kept running in the background during the inference process. All data are collected on the Xiaomi Pro using llama.cpp. Given that the additional CPU utilization introduced by the music application is below , we limit the number of parallel threads to and when running llama.cpp.
In Figure 4, for the prefill stage, background tasks cause only a minor performance drop, with differences remaining below . A similar effect is observed during the decode stage when inference is executed with threads. However, when all CPU cores are utilized (i.e., threads), decode throughput degrades more significantly under background load, with performance losses exceeding . Notably, while the background music application leaves CPU utilization largely unaffected, it reduces available RAM by several hundred megabytes, which likely contributes to the observed slowdown. These results suggest that inference performance will degrade further as the number or intensity of concurrent background tasks increases.
IV-D Parallel with AI Tasks
In addition to small workloads, there are scenarios where LLM inference must be executed alongside tasks that heavily consume compute or memory resources. For example, some image processing neural network models are deployed on mobile devices. It is important to evaluate whether the performance remains usable when these tasks run concurrently.
The experimental configuration is the same as in Section 1 except that a YOLO-based application is launched after rebooting the device and kept running in the foreground during inference. During LLM inference, the YOLO-based application activates the camera and executes the YOLOvn model for object detection. All experiments are conducted on the Xiaomi Pro with llama.cpp.
Figure 5 shows that when executed concurrently with other AI tasks, inference performance declines progressively as the number of parallel threads in the LLM inference process increases. Without additional parallel tasks, the best performance is achieved at threads, where CPU utilization reaches approximately . However, when the YOLO-based application is running (introducing an additional CPU utilization), the optimal configuration shifts to threads, which yields the smallest relative degradation. Moreover, performance degradation in the decode stage becomes significantly more pronounced than in the prefill stage. The most critical case occurs at threads: prefill throughput drops by nearly , while decode throughput collapses by more than . In practice, this indicates that during the decode stage, inference with threads (equal to the number of CPU cores) becomes essentially unusable when YOLOvn runs concurrently.
During execution with YOLOvn, the battery temperature rose from to over five inference runs, which is substantially higher compared to LLM-only execution. This induces severe thermal throttling, forcing the prime core to operate at an extremely low frequency. Consequently, optimizing processor frequency to balance performance and energy consumption is crucial for concurrent execution.
Implication 4
With YOLOv11n consuming over of CPU resources, the highest inference performance is observed with fewer parallel threads. This suggests that the optimal thread count for practical scenarios is not fixed but instead depends on real-time CPU utilization. For AI workloads that demand substantial CPU capacity, parallel execution with LLMs presents significant challenges.
Implication 5
As the number of threads executing inference increases, the throughput of the decode stage experiences a pronounced decline, indicating that the memory bottleneck is the primary limitation when LLM inference is executed concurrently with other AI workloads. Under such multi-task scenarios, intensified memory read/write pressure and bandwidth contention significantly degrade inference performance, in some cases reducing it to the point of near unusability.
V Performance: Developers’ Perspectives
@@ V-A CPU Utilization for LLM inference
V-A1 Multi-threads on CPU Cores
@@
Most popular mobile SoCs utilize the “big.LITTLE” architecture for their CPUs, which balances performance with power efficiency. This configuration typically includes multiple cores organized into two distinct clusters: “big” (prime and performance cores) and “little” (efficiency cores), as illustrated in Table 3. The detailed CPU Specifications of Apple M1 are not publicly disclosed; it is known to comprise 4 performance cores and 4 efficiency cores. While it is commonly assumed that high-load tasks are best handled by the “big” cores for optimal performance, our tests reveal that the ideal core configuration can vary across the two stages of LLM inference.
@@
Figure 6 illustrates how these core configurations impact inference speed across the four tested devices. During the prefill stage, prompt processing speed is primarily driven by the big cores. The contribution of the little cores to acceleration depends on the performance gap between the big and little cores. On devices with limited computational power, additional efficiency cores can help accelerate inference. For instance, utilizing all cores delivers the best performance on the Kirin 9000E. However, for devices with more powerful big cores, adding little cores can actually reduce performance. For example, on the Snapdragon 8 Gen 3, incorporating two little cores results in a slowdown in inference speed. This highlights the importance of optimizing core configurations based on each device’s specific CPU capabilities to maximize efficiency. Interestingly, the Dimensity 9300 benefits from using all cores because of its All-Big-Core architecture [41]. On the Apple M1, we observe that even a single performance core achieves a prefill speed comparable to that of threads. Increasing the thread count does not improve CPU utilization; instead, it may introduce excessive context switching, where threads are frequently interrupted, leaving substantial idle intervals on CPUs and reducing CPU utilization.
In the decoding phase, adding more little cores typically leads to a decline in performance. For instance, both the Kirin 9000E (Huawei Matepad12.6 Pro) and Snapdragon 8+ Gen 1 (Xiaomi Pad6 Pro) achieve their peak decoding speeds with four active big cores. This indicates that the maximum decoding speed on a device is largely dictated by the performance of the big cores. Unlike the prefill stage, decoding is more heavily constrained by memory bandwidth. This is evident in the fact that adding more big cores results in a smaller performance improvement compared to prefill, while incorporating little cores leads to a more noticeable performance drop.
Implication 6
To achieve optimal inference performance on multi-core CPUs, the decision to activate a core for inference should depend on its computational capacity relative to other cores. In general, the primary cores and performance cores contribute most to overall performance, while the efficiency cores have a negligible or even negative impact due to their limited computing resources. One potential method is distributing the workload to each core based on its compute capability and memory access overhead.
Implication 7
Since the prefill stage is compute-bound and the decoding stage is memory-bound, the core activation strategy should be adapted for each stage. Although scaling to more cores improves theoretical computation capacity, it risks introducing more synchronization and memory overhead. Take Kirin 9000E and Snapdragon 8+ Gen 1 as examples, the efficiency cores provide slight improvement during prefill, but degrade performance in decoding.
V-A2 Speedup with Special Machine Instructions
@@
Implication 8
Software developers must consider the instruction set of the target platform and optimize the code. It can involve rewriting kernel functions, adjusting the memory layout, and tailoring the computation operations, which contributes to improving performance using specialized hardware instructions. On the other hand, hardware developers should focus on designing more efficient instructions to enhance the parallelism of matrix computations and improve memory utilization.
V-B GPU Utilization for LLM inference
We note an intriguing exception regarding performance on Mali GPUs. The Mali-G Immortalis MP12, found in the Dimensity 9300 (Vivo Pad3 Pro), has better hardware parameters but has poorer performance than Adreno in Snapdragon SoC (Xiaomi14 Pro). We measure the maximum throughput for FP16 operations on devices using clpeak[45], which is a benchmarking tool intended for calculating GPU performance by executing different OpenCL kernels. The actual maximum memory bandwidth is also obtained using clpeak and presented with the theoretical maximum bandwidth in Table IV. From Table IV, we can see that the Mali-G720 MP12 exhibits times the throughput of the Adreno in float16 operations. However, Figure 3 shows that its prefill speed is times slower.
@@ V-B GPU Utilization for LLM inference
The possible reason for this discrepancy comes from the GPU utilization. Figure 8 presents the average utilization and memory bandwidth during inference on Mali-G, Adreno , and iPad Pro. It should be clarified that, on Apple devices, Load-Store utilization is reported separately as buffer reads and buffer writes, whereas on non-Apple devices we refer to the aggregate Load-Store utilization.The results reveal that MLC LLM fails to effectively utilize ALU units on both Mali and Adreno GPUs, with inefficiency being particularly severe on Mali.
During the compute-bound prefill stage, the Mali-G720’s arithmetic unit utilization averages less than . The Adreno 750 performs better, but only reaches around . By contrast, the Apple M1 achieves up to . In the memory-bound decoding stage, the Mali-G shows a severe memory bottleneck for the nearly saturated Load-Store units utilization. Taken together, these findings highlight that generic operator implementations are mismatched with the Mali GPU, leaving compute resources underused in the prefill stage and memory resources overwhelmed in the decode stage.
V-B1 Understand the underuse of non-Apple GPUs
To further investigate the causes of GPU underutilization and explore optimization opportunities, we conduct additional experiments and analyze a broader set of metrics on the Adreno and Mali-G. The results consistently highlight that the primary factors contributing to underutilization are the inefficient memory access pattern and the lack of operator optimizations tailored to specific GPU architectures.
During the prefill stage on the Adreno , cache profiling reveals L miss rates exceeding , indicating frequent cache failures and heavy reliance on system memory. Prolonged L stalls from memory transfers substantially reduce ALU activity. On the Mali-G, a similar problem arises: over of Load-Store units reads are partial-width, resulting in numerous short memory accesses that underutilize cache capacity [46]. Moreover, the rate of all registers warp reaches nearly , which halves the peak thread occupancy of the shader core and makes shader performance particularly sensitive to cache misses and memory stalls[47]. Taken together, the findings demonstrate that, although the prefill stage is inherently compute-intensive, suboptimal memory access patterns constitute the primary performance bottleneck in existing implementations.
As for the decode stage, the memory-bound nature of operators (i.e., matrix–vector multiplication) leads to frequent memory accesses for matrix operands. On the Adreno , this is reflected by an L cache miss rate approaching . The issue is even more pronounced on the Mali-G, where the average load-store unit utilization exceeds , with nearly all read operations being partial-width. Moreover, more than of external memory accesses exhibit high latency (over cycles), suggesting that the memory system is struggling to keep up with the requested data volume [48]. These findings suggest that optimizing data access patterns to reduce partial-width reads and alleviating bandwidth pressure are critical to mitigating memory bottlenecks in the decode stage.
V-B2 GPU Optimization Practices
Qualcomm’s optimized version of MLC LLM demonstrates that careful data layout design can substantially enhance GPU efficiency [49]. By reorganizing the input matrix into tiles with contiguous memory addresses, the implementation achieves improved memory coalescing. Profiling further shows that cache misses in L and L occur only at limited points (e.g., within the first ms of a ms execution). As a result, the prefill speed reaches over tokens/s, representing a six-fold improvement over the general version on the same device. This case shows that optimizing memory access patterns alone can yield substantial prefill performance gains.
Additionally, the absence of customized operator implementations further limits GPU utilization. In MLC-LLM, all operators are cross-compiled uniformly for mobile devices, resulting in identical parameter configurations for both Adreno and Mali GPUs. However, since these GPUs differ in core counts, cache capacities, and memory bandwidth, parameters of OpenCL kernels such as global work size should be tuned and optimized individually through empirical testing.
To verify the impact of kernel parameter, we evaluate different prefill chunk sizes using llama.cpp on the Xiaomi Pro. Specifically, the OpenCL kernels in llama.cpp use a fixed local work size [50], while the global block size is determined by the prefill chunk size. Large blocks typically cause memory contention, whereas small blocks underutilize compute units. Figure 9 illustrates the impact of chunk size on throughput for different prompt lengths. On the Xiaomi 14 Pro, prompts longer than 256 tokens achieve peak performance at a chunk size of 256, while shorter prompts perform best when the chunk size matches the prompt length. These optimal values, however, may vary across devices, suggesting that pre-testing or profiling is necessary to determine device-specific configurations.
Implication 9
General-purpose LLM implementations on mobile GPUs do not fully harness the parallelism capabilities of these devices. The results highlight a significant performance gap between different GPU architectures (e.g., Mali vs. Adreno)—for instance, Adreno GPU may achieve throughput than Mali GPU (which has higher computation capacity) for the same workload. To unlock the full potential of mobile GPUs, customized implementations are essential.
Implication 10
The primary performance bottleneck on GPUs arises from suboptimal data layouts and memory access patterns. Even in the compute-bound prefill stage, frequent external memory accesses cause significant ALU idle time. Optimizations such as restructuring data layouts or applying memory coalescing can improve cache utilization, while rebalancing thread workloads and minimizing register usage can further enhance computational efficiency.
V-C Impact of DVFS and Thermal Throttling
V-C1 Impact of DVFS on CPU